# ChineseSimpleVQA - “See the World, Discover Knowledge”: A Chinese Factuality Evaluation for Large Vision Language Models

### v2

@@ Abstract
The evaluation of factual accuracy in large vision language models (LVLMs) has lagged behind their rapid development, making it challenging to fully reflect these models’ knowledge capacity and reliability. In this paper, we introduce the first factuality-based visual question-answering benchmark in Chinese, named ChineseSimpleVQA, aimed at assessing the visual factuality of LVLMs across 8 major topics and 56 subtopics. The key features of this benchmark include a focus on the Chinese language, diverse knowledge types, a multi-hop question construction, high-quality data, static consistency, and easy-to-evaluate through short answers. Moreover, we contribute a rigorous data construction pipeline and decouple the visual factuality into two parts: seeing the world (i.e., object recognition) and discovering knowledge. This decoupling allows us to analyze the capability boundaries and execution mechanisms of LVLMs. Subsequently, we evaluate 34 advanced open-source and closed-source models, revealing critical performance gaps within this field.
@@ 56 LLM-as-a-Judge
Subsequently, we conducted a comprehensive evaluation of 34 advanced LVLMs, covering 8 major topics and 56 sub-topics. Here are some insightful findings:
@@ 3.5 Dataset Statistics
LLaVA-onevison-0.6B 7.9 82.3 9.8 8.8 8.3 8.5 70.6 20.8 10.8 9.5
@@ 4.2 Baseline Models
We evaluate 13 closed-source LVLMs (i.e., o1-preview , o1-mini , GPT-4o , GPT-4o-mini  [^4] https://openai.com/index/ , Gemini-1.5-pro-flash, Gemini-1.5-pro Team et al. (2024), Gemini-2.0-pro-flash [^5] https://deepmind.google/technologies/gemini/flash/, Claude-3.5-sonnet2666https://www.anthropic.com/news/claude-3-5-sonnet, Claude-3.5-sonnet, Doubao-1.5-vision-pro [^7] https://team.doubao.com/en/direction/vision, Doubao-vision-pro/lite, Qwen-VL-max Bai et al. (2023)) and 21 open-source LVLMs (i.e., Qwen2-VL series Wang et al. (2024), Qwen2.5-VL series [^8] https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api,Deepseek-VL-2 series Wu et al. (2024), LLaVA-one-vision-0.6B/72B Li et al. (2024a), GLM-4V [^9] https://open.bigmodel.cn/dev/api/normal-model/glm-4v, Llama-vision-11B/90B [^10] https://ollama.com/library/llama3.2-vision, InternVL series Chen et al. (2024).).
@@ 4.4 Main Results
  • o1-preview is the best performer for both recognition and final questions among all models. The next best are Gemini-2.0-pro-flash and Gemini-1.5-pro, which follow closely behind. For Chinese ability, Qwen family performs the best.

### v3

@@ Abstract
The evaluation of factual accuracy in large vision language models (LVLMs) has lagged behind their rapid development, making it challenging to fully reflect these models’ knowledge capacity and reliability. In this paper, we introduce the first factuality-based visual question-answering benchmark in Chinese, named ChineseSimpleVQA, aimed at assessing the visual factuality of LVLMs across 8 major topics and 56 subtopics. The key features of this benchmark include a focus on the Chinese language, diverse knowledge types, a multi-hop question construction, high-quality data, static consistency, and easy-to-evaluate through short answers. Moreover, we contribute a rigorous data construction pipeline and decouple the visual factuality into two parts: seeing the world (i.e., object recognition) and discovering knowledge. This decoupling allows us to analyze the capability boundaries and execution mechanisms of LVLMs. Subsequently, we evaluate 35 advanced open-source and closed-source models, revealing critical performance gaps within this field.
@@ 56 LLM-as-a-Judge
Subsequently, we conducted a comprehensive evaluation of 35 advanced LVLMs, covering 8 major topics and 56 sub-topics. Here are some insightful findings:
@@ 3.5 Dataset Statistics
LLaVA-onevison-7B 18.0 57.7 24.3 23.8 20.5 14.5 61.2 24.4 19.1 16.5
LLaVA-onevison-0.5B 7.9 82.3 9.8 8.8 8.3 8.5 70.6 20.8 10.8 9.5
@@ 4.2 Baseline Models
We evaluate 13 closed-source LVLMs (i.e., o1-preview (0901) , o1-mini (0901) , GPT-4o (0806) , GPT-4o-mini (0708)  [^4] https://openai.com/index/ , Gemini-1.5-pro-flash, Gemini-1.5-pro Team et al. (2024), Gemini-2.0-flash [^5] https://deepmind.google/technologies/gemini/flash/, Claude-3.5-sonnet2666https://www.anthropic.com/news/claude-3-5-sonnet, Claude-3.5-sonnet, Doubao-1.5-vision-pro [^7] https://team.doubao.com/en/direction/vision, Doubao-vision-pro/lite, Qwen-VL-max Bai et al. (2023)) and 22 open-source LVLMs (i.e., Qwen2-VL series Wang et al. (2024), Qwen2.5-VL series [^8] https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api,Deepseek-VL-2 series Wu et al. (2024), LLaVA-one-vision-0.6B/72B Li et al. (2024a), GLM-4V-9B [^9] https://open.bigmodel.cn/dev/api/normal-model/glm-4v, Llama-vision-11B/90B [^10] https://ollama.com/library/llama3.2-vision, InternVL series Chen et al. (2024).).
@@ 4.4 Main Results
  • o1-preview is the best performer for both recognition and final questions among all models. The next best are Gemini-2.0-flash and Gemini-1.5-pro, which follow closely behind. For Chinese ability, Qwen family performs the best.